OpenStudy (anonymous):

How many terms does the binomial expansion of (x^2 +2y^3)^20 contain? Please explain how to find this answer(:

OpenStudy (anonymous):

it will have 21 terms... for (x+y)^n there are n+1 terms

OpenStudy (anonymous):

wait how did u figure out it had 21 terms?

OpenStudy (anonymous):

power ie. n is 20 so n+1 = 20+1=21